it take one pipe15 hours to empty a water tank while the second pipe requires 18 hours. if the pipe is opened for 7 hours and then closed when the second pipe is opened,how long will it take the second pipe to empty the water tank pls email it to me johnreygayaso@yahoo.com
Answer by mananth(16946)

Emptying the pool is one job
Pipe I 15 hours
It does 1/15 of the job in 1 hour
Pipe II 18 hours
It does 1/18 of thejob in 1 hour

In 7 hours Pipe I will do 7/15 of the job
Balance job = 1 - 7/15 = 8/15
Pipe II does 1/18 job in 1 hour
To do 7/15 of the job it will take 7/15 / 1/18 * 1

= 8.4 hours